OpenStudy (anonymous):

find the value of x so the rectangles have so the rectangles have the same area. There are two rectangles. One rectangle has x and 12. The other rectangle has 16 and x-2. I don't know If this makes any sense but I really need help

OpenStudy (anonymous):

x*12=16(x-2) 12x=16x-32 4x=32 x=8

OpenStudy (anonymous):

are of rect = length * breadth since areas of both rect is same x * 12 = 16*(x-2)

OpenStudy (anonymous):

12x = 16x - 32 32 = 16x - 12x 32 = 4x 32/4 = x 8 = x

OpenStudy (anonymous):

thanks so much :D